If the value of land in an area is increasing 9 percent a year, how long will it take for property values to double? and if its 5.5 percent and 2.75 what would be the answers?
Answer by stanbon(75887)

If the value of land in an area is increasing 9 percent a year, how long will it take for property values to double? and if its 5.5 percent and 2.75 what would be the answers?
------
Assuming the value is compounded you get:
A(t) = Ao*(1+(r/n))^(nt)
----
Solve (1+(r/n))^(nt) = 2 for each of the rates
-----
For 9%::: (1+ 0.09)^t = 2
t = ln(2)/ln(1.09) = 8 years when rounded down.
------------------
For5.5%:: (1 + 0.055/1)^(1*t) = 2
-----
1.055)^t = 2
t = ln(2)/ln(1.055) = 13 years when rounded up
-----
For 2.75%::(1.0275)^t = 2
t = ln(2)/ln(1.0275) = 26 years when rounded up
-----
Cheers,
Stan H.
==============